### v1

# ADMM Based Semi-Structured Pattern Pruning Framework For Transformer

Abstract.
Abstract:NLP(natural language processsing) has achieved great success through the transformer model.However, the model has hundreds of millions or billions parameters,which is huge burden for its deployment on personal computer or small scale of server.To deal with it, we either make the model’s weight matrix relatively sparser, or compress attention layer. Pattern pruning ,one of the most important pruning methods, permits selecting fixed number of parameters in each divided pattern block and prunes it. However, the effect of pattern pruning is strictly limited by the sparsity within a region of weights in each layer. In this paper,we first introduced Alternating Direction Method of Multipliers(ADMM) based pattern pruning framework to reshape the distribution of activation map. Specifically, we propose to formulate the pattern pruning on transformer as a constrained optimization and use ADMM to optimize the problem. In this way, the initial dense feature maps is transformed to rather regionally sparsified ones.Therefore, we can then achieve higher compression ratio with better performance based on pattern pruning method. Additionally, this paper provides a theoretical derivations of the ADMM with local sparsity. Finally, we also extend the proposed ADMM based framework on quantization to demonstrate its generalization and use SR-STE to avoid gradient vanishing problem. We conduct extensive experiments on classification tasks over GLUE datasets. Significantly, we achieve 50% percent compression ratio while maintaining 55.4% Matthews correlation on COLA, 68.8% accuracy on RTE and overall score 80.1. Our framework also perform well on other tasks on GLUE datasets.
model compression, pruning, software, quantization, ADMM
1. Introduction
The large model,especially language large model(LLM) has archived great success for its ability to learn from large amount of data and trained on large-scale GPU clusters efficiently. Its outstanding performance is due to the proposition of Transformers(transformer, ).Transformers is kind of structure that only uses multi-head attention to process the patched data and thus greatly improve the volume of throughput. LLM can finish series of downstream tasks with a relatively universal pre-trained model and simple fine-tuning. Therefore, LLM is expected to be applied in different devices and scenarios. However, it faces the critical problems because of its complex model structure and great scale of parameters.
To ensure the performance, LLM models usually maintains great numbers of parameters. Even relative small models(BERT, ; lottery, ) consist of hundreds of millions of parameters while bigger ones’ boost to billions. These LLMs are hard to be deployed on the embedding device or mobile device while their inference speed is very slow. These challenges motivate researchers to find more effective techniques to compress(deep-compressoin, ) and designed accelerators, like AWQ(awq, ), to accelerate these models for these memory-bounded devices. These techniques include model quantization(adaptive-quantization, ), low rank decomposition(lowrank, ), knowledge distillation(knowledgedistilling, ), model sparsification and many others. They’re often combined to compress model parameters (dis-and-quantizaiton, ). This paper designs a better framework from the perspective of pruning to achieve improved compression effects.
Pruning is one of the great solutions of model sparsification, which can be categorized to unstructured pruning(cNN-pruning-with-reduction, ; iterative-pruning, ; Neural-net-pruning-why-and-how, ),, structured pruning(strucurepruning, ; structured-2, ; struturepruning-llm, ) and semi-structured pruning(NxMTransformer, ; dnn_patternpruning, ; n:M-sparsity, ; PP-Transformer, ). The basic theory of pruning is remove the parameter which is around zero meaning low importance. Unstructured pruning directly removes these individual parameters, which usually maintains the best accuracy but is unfriendly to the hardware implementation because of its irregularity. Structured pruning(strucurepruning, ) remove parameters in groups(e.g, two rows or columns), which significantly ease the hardware burden but may reduce the accuracy. Semi-structured pruning maintains the accuracy and the sparsity of specified granularity, reaching a good balance in terms of flexibility and efficiency.
We take our eyes on the pattern pruning method(dnn_patternpruning, ), a semi-structured pruning method, and try to improve it to accelerate the Transformer model. It cuts the weight matrix into several blocks. In each block, the parameters will be first removed through unstructured pruning(e.g., a 16×16 matrix can divided into four 4*4 size ones).After clustering these sparse matrix and generating the pattern pool, original dense matrix will search the best matched sparse matrix in the pool and dot product with it to get the final pruned matrix. The whole process is showed in Fig.1. This kind of pruning method simultaneously achieves the high sparseness, hardware-friendliness and good accuracy. Remarkably, pattern pruning can degenerate to the N×M pruning(NxMTransformer, ; n:M-sparsity, ; SR-STE, ). However, directly applying pattern pruning on Transformers meets many problems including low training speed and accuracy decline with high sparseness for its bigger size of weight and attention matrix than CNN. Therefore, we think it’s worth more optimization.
We gain the motivation to choose ADMM to optimize pattern pruning is for the weight matrix of Bert(and other transformer models) are with poor pattern sparsity, which refers to the submatrixs divided by pattern pruning are unbalanced for distribution on the aspect of sparsity. Thus, the effect of ADMM can alleviate the phenomenon and change the polarized distribution in this pre-pruning step into the shape that frankly spread around 50% percent of sparsity while maintaining the accuracy, as the Fig.2. shows. That’s the core of ADMM application to boost the efficiency of pruning. Meanwhile, we can gain better effect and accuracy because ADMM can also be applied to quantization jointly.
In this paper, we present a systematic ADMM(admmref, ) based pattern pruning Transformer framework with the extension of weight quantization. We make the following contributions:(1)To our best knowledge, it’s the first time applying the pattern pruning on both FFN and Attention layer to achieve the bigger compression ratio. We first introduce Sparse-refined straight-through estimator (SR-STE(SR-STE, )) in Transformer pattern pruning training process to deal with the vanishing gradient and accuracy degradation problem.(2)We first respectively formulate the pattern pruning problem and model quantization as a constrained nonconvex problem. We adopt the Alternating Direction Method of Multipliers(ADMM)(admmref, ) to systematically solve pattern pruning optimization problem on Transformer. By using ADMM, the optimization problem is then decomposed into two smaller sub problems which can be solved iteratively. In the pattern pruning problem and quantization problem, one of their sub problems can be solved by stochastic gradient descent and the other can be solved through closed-form solution derived from the equation. (3)With extensive experiments on GLUE dataset, we make detailed analysis of the effect of our framework powered by ADMM and SR-STE are efficient to optimize pattern pruning and basic quantization on whole Transformer model. These optimization methods are complimentary to other compression method including knowledge distillation(knowledgedistilling, ).
Figure 1. An illustration of pattern pruning applied in Transformer model
2. Related works
Due to the large scale of Transformer models, there is a growing interest in compressing these models to reduce their parameter size. Some studies(analyze_redundancy, ) have shown that the Transformer model has been found many redundant heads with embedded weights, which can be effectively pruned with minimal impact on accuracy. Pruning(dnn_patternpruning, ) represents a prevalent approach among the various strategies being explored for model compression. Non-structure pruning proposed by Song Han(deep-compressoin, ),energy-aware pruning(energy-aware, ) and a series of pruning methods have been proposed. Particularly, pattern pruning is a kind of semi-structured pruning method, which has also been used in DNN (dnn_patternpruning, ). PatDNN first introduce the complete pattern pruning method(including different kinds of optimization) and apply it on DNN workloads. Its theory is akin to N:M Transformer(NxMTransformer, ): divide the weight matrix into sub-matrices of a specified size. Within each sub-matrix, unstructured pruning of fixed sparsity is then applied. However, N:M pruning possesses its own unique features: N:M pruning specifically targets the removal of M contiguous parameters within a sequence of N parameters, typically located in a single row or column. In contrast, the divisions of common pattern pruning are more flexible, which divides the weight matrix into the given size visually (e.g.,3*3). The N:M Transformer made the attempt to apply the ADMM to the NLP(BERT) model and showed good effect on pruning of ADMM. From the performance and resources respect, we learn from N:M Transformer that ADMM can be widely applied to other pruning method on NLP and Transformer model.
The PP-Transformer (PP-Transformer, ) innovatively incorporates pattern pruning into the Transformer Model and implements it on FPGA hardware. However, it currently focuses only on applying pattern pruning to the dense weights of MLP layers and lacks optimization, leaving room for further improvement. Our work extends the PP-Transformer and PatDnn methods by utilizing ADMM and SR-STE compression techniques. We specifically optimize pattern pruning and apply it to the attention layer at the algorithm level, improving the efficiency of the Transformer model.
Unlike the N:M Transformer that leverages sparse N×M NVIDIA tensor cores for accelerated computations across different dimensions, our framework utilizes ADMM and other techniques to achieve competitive performance. We have successfully developed a custom FPGA hardware accelerator specifically designed to optimize pattern pruning and other techniques. This accelerator greatly improves the hardware-friendliness of the entire transformer block, including the attention mechanism. Furthermore, models incorporating our framework, such as PP-Transformer, can be utilized with TensorRT (tensorrt, ). TensorRT includes optimized operators that effectively leverage hardware features, maximizing performance gains. Similar tools like Flopoco (flopoco, ) can efficiently generate the entire Transformer model and the corresponding RTL code of operators. The availability of such tools demonstrates that the hardware implementation of our ADMM-based pattern pruning Transformer framework is both feasible and relatively straightforward.
3. Methodology
In this section, we firstly provide an introduction to the background of ADMM for neural networks, establishing the foundation for our approach. Then, we outline how we formulate the problem, clearly defining the objectives and constraints of our study. Next, this paper presents overall methodology, explaining how we leverage ADMM to develop a comprehensive framework for pattern pruning in Transformer models. Furthermore, we delve into the specific application of pattern pruning to the attention mechanism. Finally, we provide a detailed overview of the training process, describing the steps involved in training the pattern-pruned Transformer model using our proposed framework.
3.1. Background of ADMM for Neural Networks
Alternating Direction Method of Multipliers(ADMM) is a universal optimization method. It is primarily to solve the minimization of the objective function with multiple variables, which have constraints among each other. Its biggest advantage is to fully use the objective function’s decomposability to iteratively optimize several variables. ADMM simultaneously benefits from Dual Ascent and Method of Multipliers. When facing the large-scale problem or specific problem, it can decompose the overall problem into smaller subproblems, solve them in parallel and aggregate them to get the complete solution. Therefore,it was once be used to solve partial differential equations. Now it has been widely used in neural networks(DNN-ADMM-Framework, ; ADMM-NN, ; dnn_patternpruning, ) primarily for helping compress the convolution neural networks for its powerful ability to cope with the optimization of loss function.
3.2. Problem definition
Consider a Bert model with 12 transformer layers, which include muti-head attention and Feed Forward Network(FFN). Our target is to compress all FC(fully-connected) layers. Collect the weight matrix in FFN as Wi and the bias of weight as bi, where i is the layer number. The loss function of the Bert model is denoted by $f\big{(}\{\mathbf{W}_{i}\}_{i=1}^{N},\{\mathbf{b}_{i}\}_{i=1}^{N}\big{)}$. $\mathbf{N}$ represents the total number of layers.
Universally, We use ADMM to solve the this optimization problem (on most Neural Network) like:
$\underset{\{\mathrm{W}_{i}\},\{\mathrm{b}_{i}\}}{\operatorname*{minimize}}\quad f\big{(}\{\mathbf{W}_{i}\}_{i=1}^{N},\{\mathbf{b}_{i}\}_{i=1}^{N}\big{)},\quad\mathrm{subject~{}to}\quad\mathbf{W}_{i}\in\mathbf{S}_{i},i=1,\ldots,N$
Due to the diversity of the choice of $\mathbf{S}_{i}$, ADMM can be applied to the pattern pruning. For the pattern pruning, the choice of $\mathbf{S}_{i}$ is a set that contains several small matrices derived from the decomposition of larger matrices. Each small matrix in the set has a fixed sparsity. The sparsity can be manually set and changed.
3.3. Framework based on ADMM
The above problem is nonconvex and is impossible to optimize it only with the SGD (stochastic gradient descent). However, by using the framework based on ADMM, this problem can standardly be solved in an alternating optimization approach.
In this way, we have to first introduce a specific indicator function:
$\left.g_{i}(\mathbf{W}_{i})=\left\{\begin{matrix}0&\text{if W}_{i}\in\mathbf{S}_{i}\\ \infty&\text{otherwise}\end{matrix}\right.\right.$
We then introduce the auxiliary variable $\mathbf{Z}_{i}$ and rewrite the problem as
$\underset{\{\mathrm{W}_{i}\},\{\mathrm{b}_{i}\}}{\operatorname*{minimize}}\quad f\big{(}\{\mathbf{W}_{i}\}_{i=1}^{N},\{\mathbf{b}_{i}\}_{i=1}^{N}\big{)}+\sum_{i=1}^{N}g_{i}(\mathbf{Z}_{i}),\quad\mathrm{subject~{}to}\quad\mathbf{W}_{i}=\mathbf{Z}_{i},i=1,\ldots,N$
Then through the application of augmented Lagrangian method, the whole problem can be decomposed into two sub problems.
The first problem is:
$\underset{\{\mathrm{W}_{i}\},\{\mathrm{b}_{i}\}}{\operatorname*{minimize}}\quad f\big{(}\{\mathbf{W}_{i}\}_{i=1}^{N},\{\mathbf{b}_{i}\}_{i=1}^{N}\big{)}+\sum_{i=1}^{N}\frac{\rho_{i}}{2}\|\mathbf{W}_{i}-\mathbf{Z}_{i}^{k}+\mathbf{U}_{i}^{k}\|_{F}^{2}$
The $\rho_{i}$ is the hyper-parameter. The $\mathbf{U}_{i}^{k}$ is the dual variable updated in every iteration and the k in $\mathbf{U}_{i}^{k}$ presents the $\mathbf{k}_{th}$ iteration. Its update formulation is: $\mathbf{U}_{i}^{k}:=\mathbf{U}_{i}^{k-1}+\mathbf{W}_{i}^{k}-\mathbf{Z}_{i}^{k}$. This problem becomes convex and can be solved by stochastic gradient descent. We only need to add the extra loss $\sum_{i=1}^{N}\frac{\rho_{i}}{2}\|\mathbf{W}_{i}^{k+1}-\mathbf{Z}_{i}+\mathbf{U}_{i}^{k}\|_{F}^{2}$ generated by ADMM, and train the model normally.
The second problem is:
$\underset{\{\mathbf{Z}_{i}\}}{\operatorname*{minimize}}\sum_{i=1}^{N}g_{i}(\mathbf{Z}_{i})+\sum_{i=1}^{N}\frac{\rho_{i}}{2}\|\mathbf{W}_{i}^{k+1}-\mathbf{Z}_{i}+\mathbf{U}_{i}^{k}\|_{F}^{2}$
Though this problem is convex, but it can be smartly dealt by applying $\mathbf{Z}_{i}$ with Euclidean project onto the set $\mathbf{S}_{i}$.
$\mathbf{Z}_{i}^{k+1}=\Pi_{\mathbf{S}_{i}}(\mathbf{W}_{i}^{k+1}+\mathbf{U}_{i}^{k})$
It can trivially be seen that it makes $g_{i}(\mathbf{Z}_{i})$ equal to zero while maintains the distance between $\mathbf{Z}_{i}$ and $\mathbf{W}_{i}^{k+1}+\mathbf{U}_{i}^{k}$ as short as possible.
Based on the different problems, the Euclidean projection are variant. For the single pattern pruning problem, $\mathbf{Z}_{i}^{k+1}$ equals to the result matrix of $\mathbf{W}_{i}^{k+1}+\mathbf{U}_{i}^{k}$ after pattern pruning. That is to say, The matrices obtained by dividing a matrix exhibit the same sparsity. For the single quantization problem, $\mathbf{Z}_{i}^{k+1}$ equals to the result of setting all the values in $\mathbf{W}_{i}^{k+1}+\mathbf{U}_{i}^{k}$ as quantization values. For the ADMM optimization on pattern pruning, we do several experiments and compare the results between a direct implementation and the results obtained after employing the ADMM framework.
3.4. Pattern pruning of attention
We not only choose the dense weight matrix of the model, but also the attention mechanism to apply the pattern pruning to extend our framework to a full-stack sparsification. In attention layer, after the matrix multiplication of Query and Key, we do the pattern pruning to sparsify their results, which owns the biggest potential to be hardware-friendly. By partitioning the attention map into multiple blocks, we can easily observe the relative importance of each block based on their numerical values. As shown in Fig.2., the dense attention map is transformed into a sparse attention map through pattern pruning. The first and last blocks exhibit greater importance.
Figure 2. Pruning process of attention
Figure 3. The training process of single optimization problem
3.5. The whole training process
The whole training process is shown is Fig.3. Firstly we initialize both $\mathbf{Z}_{i}^{0}$ and $\mathbf{U}_{i}^{0}$ as zero matrix. In the first iteration, we gain the weight $\mathbf{W}_{i}^{0}$. Then, project $\mathbf{Z}_{i}^{0}$ and update $\mathbf{U}_{i}^{0}$. Next, we add the extra part generated by ADMM to the loss to deal with the first subproblem. We normally train the model and enter the next iteration. Repeat the above steps until the model converges. After ADMM optimization is done, we will keep the existing sparsity distribution. Then prune and retrain the model to fit in the downstream task.
4. Evaluation
In this section, we evaluate ADMM-based framework and show its effectiveness in compressing Transformer networks over a wide range of NLP tasks.
Implementation. We conduct experiments on NVIDIA 3090 GPU with batch size 24, and use Adamw optimizer(ADAMW, ) with the initial learning rate of 2e-5 meanwhile using learning rate scheduler to make adaptive adjustments during the training process. The code is implemented based on hugging face using python 3.9.3, pytorch 1.13.1 and cuda 12.3.
Former works have indicated that higher sparsity leads to greater performance loss when the pattern size is larger than 4*4. Therefore, for evaluation, we focus on evaluating 50% sparsity, 4*4 pattern size, 32 pattern number since it is applicable in most situations. We use pretrained model checkpoints for BERT2 , provided by the HuggingFace model repository. All models were fine-tuned. Finetuning these models required between 7 minutes (RTE) and 7 hours (MNLI) depending on the scale of the task. We further choose 4*4 pattern size with 80% sparsity and 8*8 pattern size with 50 % sparsity to demonstrate the great effect of the model weight distribution change.
Datasets. We evaluate ADMM-based pattern pruning framework and our baselines using the General Language Understanding Evaluation (GLUE) benchmark(glue, ), a collection of NLP tasks varying in data availability and complexity. We report the Spearman correlation for STS-B, the F1 score for MRPC, Matthews correlation for CoLA, and accuracy for all remaining tasks. The reported average is the geometric mean of reported scores. With GLUE as an evaluation criteria, we can better compare our framework with other pruning methods.
4.1. Main results
Bert: Baseline Bert base model
ADMM unstructured:To measure the accuracy cost of semi-structured accuracy specifically, we create another baseline that uses ADMM but induces unstructured sparsity at 50% per-layer (rather than global) sparsity.
N×M Transformer: The tranformer framework pruned with N:M method and optimized with ADMM.
Pattern pruning(without ADMM): The tranformer framework pruned with pattern pruning method with no ADMM optimization.
Pattern pruning(optimized with ADMM): The tranformer framework pruned with pattern pruning method optimized with ADMM .
We search for the optimal learning rates (7e-5) and weight decay (0.001) to achieve the best possible results. The parameter rho in the ADMM loss is adjusted in the experiment process and finally through the change process of rho, we fix rho as 0.01 and gain the good score on different tasks comprehensively. Table 1 shows the evaluation results for BERT, and the following key observations can be made. First, the basic pruning-based method sparsifies the weights of Transformer blocks but cannot explicitly satisfy the underlying hardware constraints, e.g., the pattern sparsity. While the highest accuracy on downstream tasks was preserved (81.3 vs. 81.8 on average), the obtained sparse weights had a random structure of non-zero weights, which was inefficient and unacceptable. Second, the original pattern pruning method, which was not optimized using the ADMM technique, are demonstrated to have a relatively poor performance, both in individual downstream tasks and on average(78.1%). In contrast, the pattern pruning framework optimized through ADMM demonstrated accuracy levels in various tasks that approached those of the N×M Transformer, which had been optimized by multiple methods. Notably, it achieved scores of 55.4 and 68.8 on the COLA and RTE tasks, surpassing the N×M Transformer. The overall average score reached 80.1, which was significantly higher than that of the pre-optimization pattern pruning method by near 2 percent.
Table 1. Results on GLUE benchmark
Model Task Average
MNLI(m/mm) SST-2 QNLI CoLA STS-B MRPC RTE
Samples 392k 67k 108k 8.5k 5.7k 5.5k 2.5k
Baseline 84.5/84.8 92.5 91.6 56.7 89.6 91.7 70.7 81.8
ADMM(unstructured) 84.0/84.7 92.5 91.0 57.5 89.6 90.5 68.2 81.3
N×M Transformer 82.3/83.4 92.3 90.4 55.3 89.3 90.8 68.6 80.5
pattern pruning(without ADMM) 80.6/81.5 90.9 88.8 53.7 87.3 85.1 65.3 78.1
pattern pruning(optimized with ADMM) 82.0/83.1 92.0 89.9 55.4 88.4 89.2 68.8 80.1
4.2. Effect on weight distribution
Furthermore, we investigate the impact of our ADMM-based framework on the weight distribution of each layer. Fig.4. illustrates the alteration in the weight distribution of 0th layer of Bert following the training process within our ADMM-based framework. Our framework can effectively tailor to different shape’s weight and according to the division to purposefully make sparsity of each part approach our given sparsity.Here we set it as 50%. In Fig.4., ADMM-optimized weight matrix (the query weight of layer 0) exhibits a notable increase in zero values compared to the original matrix. This demonstrates the effectiveness of the ADMM optimization in achieving a 50% sparsity distribution for pattern pruning.
What’s more, the extensive Query weight matrix (768,768) is partitioned into nine discrete sections, each with a dimension of 256. This approach is employed to illustrate the impact of the ADMM optimization framework on the distribution of weights within each section. The weight distributions of the nine independent parts are presented in Fig. 5. It can be observed that the majority of these parts exhibit a similar trend to the original matrix, approaching 50% sparsity and becoming increasingly sparse. These findings prove the effectiveness of the proposed approach in altering weight distributions and enhancing hardware compatibility.
Figure 4. Distribution change of query weight of layer 0 before/after ADMM for pattern pruning
Figure 5. Distribution of each part of the query weight
4.3. Analysis of results
Our best accuracy of Bert model after optimization, before pruning, achieves 89.7%. After fully pruning the model, the average score of Bert model can achieve 80.1% while maintaining 50% percent of sparsity. On specific task, COLA and RTE, it can outperform the N×M Transformer, which is an excellent framework nowadays.
Validation and hardware-friendness improvement Our evaluation scores and hardware efficiency are clearly demonstrated in the data and charts. The significant improvements can be attributed to our ADMM optimization framework. Within this framework, we integrated the SR-STE method to prevent gradient vanishing, enhancing the training effect.During pattern pruning, varying sparsity levels in different weight blocks led to substantial discrepancies between the pruned matrices and the original large matrices, either by excessively reducing large values or retaining too many small values in certain blocks, severely affecting validation accuracy. Our ADMM-based framework iteratively optimizes the weight distribution, ensuring that the weights are more evenly distributed to approximately 50% sparsity before pattern pruning. This approach significantly boosts the post-pruning scores, achieving an overall score of 80.1, which is 2 percentage points higher than pre-optimization and only 0.4 percentage points lower than one of the best current sparse frameworks, the N×M Transformer.Additionally, the structural uniformity achieved through our pruning method provides hardware efficiency comparable to the N×M Transformer. By leveraging and improving the existing pattern pruning FPGA hardware framework, we can achieve excellent deployment results.
5. Future works
According to PP-Transformer(PP-Transformer, ), the performance of pattern pruning(bigger pattern) will drop quickly when sparsity becomes above 80%percent. In our experiment process, we found that after ADMM improvement, the performance with high sparsity, bigger pattern choice pattern pruning significantly can partly deal with that problem, but still with unacceptable performance drop. Therefore, we want to find other optimization method and combine it with ADMM to completely stop it from dropping. On the other hand, we hope we can develop a complete set of hardware framework tailoring to ADMM optimization, which will greatly speed up the deploy and run time on FPGA or ASIC chips.
6. Conclusions
ADMM can bring the transformer weight distribution to the shape we desire and greatly improve the pruning performance. This paper demonstrate the effectiveness of ADMM on pattern pruning from performance and weight distribution aspects. Furthermore, ADMM based pattern pruning Framework is orthogonal to quantization and other optimization techniques. This framework we proposed can be wildly applied to almost all the models based on the transformer. including Bert,Vit and others.While the framework do introduce a loss and accuarcy gap because of compression, we believe the ADMM based pattern pruning framework will work and keep its importance.
7. Acknowledgement
We sincerely thank my advisor ZH.L and the anonymous reviewers for their constructive comments.